higher grain prices caused the value of Indiana cropland to rise 14 % from 1995 to 1996.An acre of farmland was sold for $3078 in 1996. How much would it have been worth in 1995 ?
Answer by ikleyn(52915)

.

x + 0.14x = 3078

1.14x     = 3078

    x     = 3078/1.14 = 2700.    ANSWER

Solved.